# Distribution of lengths of closed saddle connections on moduli space of large genus translation surface

### v2

@@ 5.1. The measure of exception set
Combine the two proposition we have
@@ 5.2. Collapsing simultaneously
Let $\mathcal{C}^{\prime}_{g,Per}=\mathcal{C}^{1,0}_{g,Per}(2^{g-1})\setminus(\mathcal{C}_{g}^{1}(\frac{B}{\sqrt{g}})\cup\mathcal{C}_{g}^{2}(\frac{B}{\sqrt{g}}))$, we will adjust the surgery from last section on $\mathcal{C}^{\prime}_{g,Per}$ to collapse the closed saddle connections in $\Gamma_{1},\cdots,\Gamma_{k}$ in order. The opening and pinching operations are same to the situation of one curve. The only modification is the equation 4.1: by remark 4.3, we need to choose the moving vector $\upsilon$ to ensure the area unchanged, hence we can choose a basis containing $\gamma^{\prime}_{1},\gamma_{1}\cdots,\gamma^{\prime}_{K},\gamma_{K}$, where $\gamma^{\prime}_{i}$ is a saddle connection from opening $\gamma_{i}$. Then the moving vector $\upsilon$ can be constructed like equation 4.1: after defining the value on $\gamma_{i}$, we can choose another relative homotopy class and give a value of $\upsilon$ on it to ensure the area unchanged. This means we open and move the closed saddle connections simultaneously and then pinch them.
@@
$\hat{F}:\hat{\mathcal{C}}^{\prime}_{g,Per}(2^{g-1})\to M_{K}\cdot\tilde{{\mathcal{H}}}_{g-K}(2^{g-1-K},0^{2K})\times\prod^{k}_{i=1}(A_{[\frac{a_{i}}{\sqrt{g}},\frac{b_{i}}{\sqrt{g}}]})^{n_{i}},$
where $\hat{\mathcal{C}}^{\prime}_{g,Per}(2^{g-1})$ is defined as $\hat{\mathcal{C}}^{1,0}_{g}(2^{g-1})$ and $M_{K}$ is a combinatorial data consisting of choosing $K$ ordered zeros from g1g-1 zeros to label. As the discussion above, we have
@@
$=\operatorname{Vol}(\hat{\mathcal{C}}^{\prime}_{g}(2^{g-1})))$
@@
Since we don’t need to deal with simple zeros in the collapsing surgery, samilarly to the proof of theorem 1.2, we can prove Corollary 1.4.
@@ 6. General stratum
This section we consider the stratum ${\mathcal{H}}_{g}(m^{O(g)},1^{2g-2-mO(g)})$, where $m\geq 3$.
If a stratum has zeros of order more than $3$, the angles at both sides are not certain. But we can consider some fixed configuration of closed saddle connection. For a configuration $(J,b^{\prime}_{k},b^{\prime\prime}_{k},a^{\prime}_{i},a^{\prime\prime}_{i})$ of closed saddle connection. We have known it suffices to consider multiplicity $1$ without cylinders, then the configuration become $(b^{\prime},b^{\prime\prime})$, where $b^{\prime}+b^{\prime\prime}=m$, which means the angles of the closed saddle connection are $(2b^{\prime}+1)\pi$ on one side and $(2b^{\prime\prime}+1)\pi$ on the other side. Now we choose all such configuration at each $m$-order zeros.
@@
$\hat{F}:\hat{\mathcal{C}}^{\prime}_{g,Per}(m^{O(g)},1^{O^{\prime}(g)})\to\tilde{{\mathcal{H}}}_{g-K}((b^{\prime}-1)^{K},(b^{\prime\prime}-1)^{K},m^{O(g)-K},1^{O^{\prime}(g)})\times\prod^{k}_{i=1}(A_{[\frac{a_{i}}{\sqrt{g}},\frac{b_{i}}{\sqrt{g}}]})^{n_{i}},$
where $O^{\prime}(g)=2g-2-mO(g)$. Again we have
$\lim_{g\to\infty}\mathbb{E}[(L^{\prime}_{g,[a_{1},b_{1}]})_{n_{1}}\cdots(L^{\prime}_{g,[a_{k},b_{k}]})_{n_{k}}]=\prod^{k}_{i=1}(\frac{(m+1)\pi(b_{i}^{2}-a_{i}^{2})}{b^{\prime}b^{\prime\prime}})^{n_{i}}.$
And similar to the proof of theorem 1.2 we have
@@
For the stratum ${\mathcal{H}}(m^{O(g)},1^{2g-2-mO(g)})$, a configuration $(b^{\prime},b^{\prime\prime})$ and $a,b\in\mathbb{R}$, let the random variable
@@
$\lambda_{[a_{i},b_{i}]}=\frac{(m+1)\pi(b_{i}^{2}-a_{i}^{2})}{b^{\prime}b^{\prime\prime}}.$

### v3

@@ 5.1. The measure of exception set
Combine the two propositions we have
@@ 5.2. Collapsing simultaneously
Let $\mathcal{C}^{\prime}_{g,Per}=\mathcal{C}^{1,0}_{g,Per}(2^{g-1})\setminus(\mathcal{C}_{g}^{1}(\frac{B}{\sqrt{g}})\cup\mathcal{C}_{g}^{2}(\frac{B}{\sqrt{g}}))$, we will adjust the surgery from last section on $\mathcal{C}^{\prime}_{g,Per}$ to collapse the closed saddle connections in $\Gamma_{1},\cdots,\Gamma_{k}$ in order. The opening and pinching operations are same to the situation of one curve. The only modification is the moving surgery: we choose the shortest homology class $\beta$ on $X\setminus\bigcup\limits_{i=1}^{k}\gamma_{i}$ which is non-separable and non-homotopic to the boundaries, and choose its dual curve $\alpha$. Then we choose the tangent vector $\upsilon$ satisfying $\upsilon(\gamma_{i})=-\omega(\gamma_{i})$, and adjust the value of $\upsilon(\alpha)$ to ensure the area unchanged. Since $k$ is fixed and $g\to\infty$, this $\upsilon$ can be chosen.
@@
$\hat{F}:\hat{\mathcal{C}}^{\prime}_{g,Per}(2^{g-1})\to M_{K}\cdot{\mathcal{H}}_{g-K}(2^{g-1-K},0^{2K})\times\prod^{k}_{i=1}(A_{[\frac{a_{i}}{\sqrt{g}},\frac{b_{i}}{\sqrt{g}}]})^{n_{i}},$
where $\hat{\mathcal{C}}^{\prime}_{g,Per}(2^{g-1})$ is defined as $\hat{\mathcal{C}}^{1,0}_{g}(2^{g-1})$ and $M_{K}$ is a combinatorial data consisting of choosing $K$ ordered zeros from g1g-1 zeros to label. From section 4.2.4, this map is one to one on a total measure subset of ${\mathcal{H}}_{g-K}(2^{g-1-K},0^{2K})$, since at a marked point there is only one choice to split along a given vector.
As the discussion above, we have
@@
$=\operatorname{Vol}(\hat{\mathcal{C}}^{\prime}_{g}(2^{g-1}))$
@@
Then by Theorem 2.5 the Theorem 1.2 is proved.
@@ 6. General stratum
This section we first consider the stratum ${\mathcal{H}}_{g}(m^{O(g)},1^{2g-2-mO(g)})$, where $m\geq 3$ and $\lim\limits_{g\to\infty}\frac{O(g)}{g}=c$. If a stratum has zeros of order more than $3$, the angles at both sides are not certain. But we can consider some fixed configuration of closed saddle connection. For a configuration $(J,b^{\prime}_{k},b^{\prime\prime}_{k},a^{\prime}_{i},a^{\prime\prime}_{i})$ of closed saddle connection. We have known it suffices to consider multiplicity $1$ without cylinders, then the configuration become $(b^{\prime},b^{\prime\prime})$, where $b^{\prime}+b^{\prime\prime}=m$, which means the angles of the closed saddle connection are $(2b^{\prime}+1)\pi$ on one side and $(2b^{\prime\prime}+1)\pi$ on the other side. Now we choose all such configuration at each $m$-order zeros.
@@
$\hat{F}:\hat{\mathcal{C}}^{\prime}_{g,Per}(m^{O(g)},1^{O^{\prime}(g)})\to M_{K}\cdot{\mathcal{H}}_{g-K}((b^{\prime}-1)^{K},(b^{\prime\prime}-1)^{K},m^{O(g)-K},1^{O^{\prime}(g)})\times\prod^{k}_{i=1}(A_{[\frac{a_{i}}{\sqrt{g}},\frac{b_{i}}{\sqrt{g}}]})^{n_{i}}.$
In this situation, the construction of the inverse mapping is a little different. Suppose we have chosen a fixed vector sequence $(v_{1},\cdots,v_{K})\in\prod^{k}_{i=1}(A_{[\frac{a_{i}}{\sqrt{g}},\frac{b_{i}}{\sqrt{g}}]})^{n_{i}}$, for some $v_{i}$, we want split the surface in ${\mathcal{H}}_{g-K}((b^{\prime}-1)^{K},(b^{\prime\prime}-1)^{K},m^{O(g)-K},1^{O^{\prime}(g)})$ along $v_{i}$ at a pair of zeros of orders $(b^{\prime}-1)$ and $(b^{\prime\prime}-1)$, but we have $b^{\prime}b^{\prime\prime}$ choices to choose the split locus. So the mapping $\hat{F}$ is $(b^{\prime}b^{\prime\prime})^{K}$ to one. So we have
$\operatorname{Vol}(\hat{\mathcal{C}}^{\prime}_{g,Per}(m^{O(g)},1^{O^{\prime}(g)}))=$
$|M_{K}|(b^{\prime}b^{\prime\prime})^{K}\operatorname{Vol}({\mathcal{H}}_{g-K}((b^{\prime}-1)^{K},(b^{\prime\prime}-1)^{K},m^{O(g)-K},1^{O^{\prime}(g)}))\prod^{k}_{i=1}(\pi\frac{b_{i}^{2}-a_{i}^{2}}{g})^{n_{i}},$
where $|M_{K}|=O(g)(O(g)-1)\cdots(O(g)-K)$. And similar to the proof of theorem 1.2 we have
$\lim_{g\to\infty}\mathbb{E}[(L^{\prime}_{g,[a_{1},b_{1}]})_{n_{1}}\cdots(L^{\prime}_{g,[a_{k},b_{k}]})_{n_{k}}]$
$=|M_{K}|(b^{\prime}b^{\prime\prime})^{K}\frac{\operatorname{Vol}({\mathcal{H}}_{g-K}((b^{\prime}-1)^{K},(b^{\prime\prime}-1)^{K},m^{O(g)-K},1^{O^{\prime}(g)}))}{\operatorname{Vol}({\mathcal{H}}_{g}(m^{O(g)},1^{O^{\prime}(g)}))}\prod^{k}_{i=1}(\pi\frac{b_{i}^{2}-a_{i}^{2}}{g})^{n_{i}}$
$=(b^{\prime}b^{\prime\prime})^{K}\frac{(m+1)^{K}}{(b^{\prime}b^{\prime\prime})^{K}}\prod^{k}_{i=1}[\pi(b_{i}^{2}-a_{i}^{2})]^{n_{i}}\frac{O(g)(O(g)-1)\cdots(O(g)-K)}{g^{K}}$
$=c^{K}(m+1)^{K}\prod^{k}_{i=1}[\pi(b_{i}^{2}-a_{i}^{2})]^{n_{i}}$
$=\prod^{k}_{i=1}[c\pi(m+1)(b_{i}^{2}-a_{i}^{2})]^{n_{i}}.$
So we have
@@
Suppose $\lim\limits_{g\to\infty}\frac{O(g)}{g}=c$. For the stratum ${\mathcal{H}}(m^{O(g)},1^{2g-2-mO(g)})$, a configuration $\mathcal{C}=(b^{\prime},b^{\prime\prime})$ and $a,b\in\mathbb{R}$, let the random variable
@@
$\lambda_{[a_{i},b_{i}]}=c(m+1)\pi(b_{i}^{2}-a_{i}^{2}).$
Note that $\lambda_{[a_{i},b_{i}]}$ is independent to $(b^{\prime},b^{\prime\prime})$, now we consider general closed saddle connections in ${\mathcal{H}}(m^{O(g)},1^{2g-2-mO(g)})$. Let
$L_{g,[a,b]}:{\mathcal{H}}(m^{O(g)},1^{2g-2-mO(g)})\to\mathbb{N}_{0}$
be the number of closed saddle connections on $(X,\omega)$ with lengths in $\frac{a}{\sqrt{g}},\frac{b}{\sqrt{g}}]$. Then we have
$L_{g,[a,b]}=\sum_{i=1}^{m-1}L_{\mathcal{C}_{i},g,[a,b]},$
where $\mathcal{C}_{i}=(i,m-i)$.
Fix some $i,j\in\{1,2,\cdots,m-1\}$, let
$\hat{L}^{i,j}_{g,[a,b]}=(L_{\mathcal{C}_{i},g,[a,b]},L_{\mathcal{C}_{j},g,[a,b]}).$
We will compute the distribution of $\hat{L}^{i,j}_{g,[a,b]}$. For $n_{i},n_{j}$, consider the set consisting of
$(X,\omega,\Gamma_{i},\Gamma_{j}),$
where $(X,\omega)\in\mathcal{C}^{1,0}_{g}(m^{O(g)},1^{2g-2-mO(g)})$, and $\Gamma_{i}$ is an ordered list of $n_{i}$ closed saddle connections $(\gamma_{1,i},\cdots,\gamma_{n_{i},i})$ with lengths in $\frac{a}{\sqrt{g}},\frac{b}{\sqrt{g}}]$ and configuration $\mathcal{C}_{i}$.
If we want to collapse curves in $\Gamma_{i}$ and $\Gamma_{j}$, we need to consider the exception set where the surgery loci intersect. Similar to 5.1, if two surgery loci intersect, then either the two zeros are close or the two closed saddle connections share the same zero. From section 5.1, the measure of such subset goes to zero as $g\to\infty$. So we can also collapse the curves in $\Gamma_{i}$ and $\Gamma_{j}$ simultaneously. Again we have
Proposition 6.2.
$\hat{L}^{i,j}_{g,[a,b]}$ converges jointly in distribution to a vector of random variables with Poisson distributions of means $\lambda_{[a,b]}$, where
$\lambda_{[a,b]}=c(m+1)\pi(b^{2}-a^{2}).$
This means for different configurations, the limits of $L_{\mathcal{C}_{i},g,[a,b]}$ are independent.
From the property of Poisson distribution, we have
Proposition 6.3.
$L_{g,[a,b]}=\sum\limits_{i=1}^{m-1}L_{\mathcal{C}_{i},g,[a,b]}$ converges in distribution to a Poisson distributions of mean $\lambda_{[a,b]}$, where
$\lambda_{[a,b]}=c(m^{2}-1)\pi(b^{2}-a^{2}).$
Proof.
Since $L_{\mathcal{C}_{i},g,[a,b]}$ are independent and
$L_{g,[a,b]}=\sum_{i=1}^{m-1}L_{\mathcal{C}_{i},g,[a,b]}.$
We have $L_{g,[a,b]}$ converges in distribution to a Poisson distribution, and the mean is $(m-1)\lambda_{[a,b]}=c(m^{2}-1)\pi(b^{2}-a^{2})$. ∎
Finally, consider ${\mathcal{H}}(m_{1}^{O_{1}(g)},\cdots,m_{k}^{O_{k}(g)},1^{O(g)})$, where $\lim\limits_{g\to\infty}\frac{O_{i}(g)}{g}=c_{i}.$ Let
$L_{g,i,[a,b]}:{\mathcal{H}}(m_{1}^{O_{1}(g)},\cdots,m_{k}^{O_{k}(g)},1^{O(g)})\to\mathbb{N}_{0}$
be the number of closed saddle connections at a zero of order $m_{i}$ on $(X,\omega)$ with lengths in $\frac{a}{\sqrt{g}},\frac{b}{\sqrt{g}}]$. From Proposition 6.3 we have $L_{g,i,[a,b]}$ converges in distribution to a Poisson distributions of mean $\lambda_{i,[a,b]}$, where
$\lambda_{i,[a,b]}=c_{i}(m_{i}^{2}-1)\pi(b^{2}-a^{2}).$
Fix some $i,j\in\{1,2,\cdots,k-1\}$, let
$\hat{L}^{i,j}_{g,[a,b]}=(L_{g,i,[a,b]},L_{g,j,[a,b]}).$
Similarly if we want to compute the distribution of $\hat{L}^{i,j}_{g,[a,b]}$, we need to consider the set
$(X,\omega,\Gamma_{i},\Gamma_{j}),$
where $(X,\omega)\in\mathcal{C}^{1,0}_{g}(m_{1}^{O_{1}(g)},\cdots,m_{k}^{O_{k}(g)},1^{O(g)})$, and $\Gamma_{i}$ is an ordered list of $n_{i}$ closed saddle connections at a zero of order $m_{i}$. Again from 5.1, these saddle connections can be collapsed simultaneously except for a measure zero set. Then we have $L_{g,i,[a,b]}$ and $L_{g,j,[a,b]}$ converge to independent Poisson distributions of means $\lambda_{i,[a,b]}$ and $\lambda_{j,[a,b]}$.
Let $L_{g,[a,b]}$ be the number of closed saddle connections on $(X,\omega)$ with lengths in $\frac{a}{\sqrt{g}},\frac{b}{\sqrt{g}}]$, where $(X,\omega)\in{\mathcal{H}}(m_{1}^{O_{1}(g)},\cdots,m_{k}^{O_{k}(g)},1^{O(g)})$. From the property of Poisson distribution, we have $L_{g,[a,b]}=\sum\limits_{i=1}^{k-1}L_{g,i,[a,b]}$ converges in distribution to a Poisson distributions of mean $\lambda_{[a,b]}$, where
$\lambda_{[a,b]}=\sum_{i=1}^{k-1}\lambda_{i,[a,b]}=\sum_{i=1}^{k-1}c_{i}(m_{i}^{2}-1)\pi(b^{2}-a^{2}).$
Then we have proved Theorem 1.4.